# CLIP-EBC: CLIP Can Count Accurately through Enhanced Blockwise Classification

### v2

@@ 3.2 The Structure of CLIP-EBC
Next, the resulting $n$ text prompts are tokenized using CLIP’s tokenizer. We then input the tokenized text into the original CLIP text encoder, with its parameters frozen during training. This process yields text embeddings $\boldsymbol{F}^{\text{txt}}$ with dimensions $d\times n$.
@@
The probability map $\boldsymbol{P}^{*}$ can be computed from the image feature map $\boldsymbol{F}^{\text{img}}$ and the text features $\boldsymbol{F}^{\text{txt}}$. First, we calculate the cosine similarity between the image feature vector $\boldsymbol{F}^{\text{img}}_{:,\,i,\,j}$ and the $n$ extracted text embeddings. We then normalize these similarities using softmax to obtain $\boldsymbol{P}^{*}_{:,\,i,\,j}$, which denotes the probability scores across the $n$ bins for block $(i,\,j)$. To obtain the predicted density map $\boldsymbol{Y}^{*}$, we use Eq. (2). We train CLIP-EBC with our loss function in Eq. (3.1).
Methods Type ShanghaiTech A ShanghaiTech B UCF-QNRF NWPU (Val)
MAE RMSE MAE RMSE MAE RMSE MAE RMSE
CrowdCLIP (Liang et al. 2023) U 146.1 236.3 69.3 85.8 283.3 488.7 - -
DMCount (Wang et al. 2020a) R 59.7 95.7 7.4 11.8 85.6 148.3 70.5 357.6
GLoss (Wan, Liu, and Chan 2021) R 61.3 95.4 7.3 11.7 84.3 147.5 79.3 346.1
P2PNet (Song et al. 2021a) R 52.7 85.0 6.3 9.9 85.3 154.5 77.4 362.0
CLTR (Liang, Xu, and Bai 2022) R 56.9 95.2 6.5 10.2 85.8 141.3 61.9 246.3
STEERER (Han et al. 2023) R 54.5 86.9 5.8 8.5 74.3 128.3 63.7 309.8
S-DCNet (Xiong et al. 2019) C 58.3 95.0 6.7 10.7 104.4 176.1 - -
UEPNet (Wang et al. 2021) C 54.6 91.2 6.4 10.9 81.1 131.6 - -
DC-Reg (Xiong and Yao 2022) C 59.8 100.0 6.8 11.5 84.8 142.3 - -
CLIP-EBC (ResNet50, ours) C 54.0 83.2 6.0 10.1 80.5 136.6 38.6 90.3
CLIP-EBC (ViT/B-16, ours) C 52.5 85.9 6.6 10.5 80.3 139.3 36.6 81.7
Table 3: Comparison of our model CLIP-EBC with recent crowd counting approaches on ShanghaiTech A & B (Zhang et al. 2016), UCF-QNRF (Idrees et al. 2018) and NWPU-Crowd (Val) (Wang et al. 2020b). The “Type” column groups all methods into unsupervised (“U”), regression-based (“R”), or classification-based (“C”). The best results are highlighted in bold, and the second best results are underlined. Our model CLIP-EBC demonstrates competitive performance across all datasets, with particularly significant superiority on the NWPU validation split, where it outperforms all existing methods.
4 Experiments
In this section, we first empirically validate the effectiveness of each enhancement strategy in our EBC framework (Sec. 4.1), followed by a comparison of our CLIP-EBC model against state-of-the-art methods on four widely used crowd-counting benchmark datasets: ShanghaiTech A and B (Zhang et al. 2016), UCF-QNRF (Idrees et al. 2018), and NWPU-Crowd (Wang et al. 2020b) (Sec.4.2). In Sec. 4.3 we analyze the impact of bin granularity, while in Sec. 4.4 we investigate the effects of varying levels of noise control. Additional experiments are provided in the Appendix.
Training Details
@@ 4 Experiments
We set the block size $r=8$ (unless explicitly specified), which is commonly used in many works (Li, Zhang, and Chen 2018; Liu, Salzmann, and Fua 2019; Wang et al. 2020a). For this block size, we set $m=4$. The CLIP-EBC models are initialized with the weights from CLIP (Radford et al. 2021). For the remaining models, we initialize the encoder with pre-trained weights on ImageNet, while the decoder (if applicable) is initialized with random values drawn from a normal distribution. In the proposed DACE loss (3.1), we set $\lambda$ to 1, unless specified otherwise. We use the Adam optimizer (Kingma and Ba 2015) to train all our models with an initial learning rate of $41e-$ 4, which is adjusted through a cosine annealing schedule (Loshchilov and Hutter 2017). The batch size is fixed at $8$ for all datasets. For our ViT-based CLIP-EBC model, we apply visual prompt tuning (Jia et al. 2022) by prepending 32 learnable tokens to each layer, rather than fine-tuning the entire backbone. All models are implemented using PyTorch 2.2.1 and trained on Ubuntu 22.04.4 LTS.
Evaluation metrics
Following existing methods, we use the mean absolute error (MAE) and root mean square error (RMSE) to evaluate our models, defined as follows:
[…]
where $N$ is the number of images in the test set, $C_{i}$ is the ground-truth global count value of image $i$, and $C_{i}^{*}$ is the predicted count value obtained by integrating over the predicted density map $\boldsymbol{Y}^{*}$, which can be obtained by Eq. (2). Notably, lower scores indicate better performance.
@@
P2PNet (Song et al. 2021a) 72.6 331.6
CLTR (Liang, Xu, and Bai 2022) 74.4 333.8
STEERER (Han et al. 2023) 63.7 309.8
CrowdHat (Wu and Yang 2023) 68.7 296.9
PET (Liu et al. 2023) 74.4 328.5
@@
Table 4: Comparison of our model CLIP-EBC with state-of-the-art methods on NWPU-Crowd (Test) (Wang et al. 2020b). CLIP-EBC achieves both the lowest MAE and RMSE on this large-scale benchmark database, significantly surpassing all strong baselines.
4.1 Effectiveness of EBC
We first demonstrate the effectiveness of our EBC framework using VGG16 (Simonyan and Zisserman 2015) as the backbone, trained and evaluated on the UCF-QNRF dataset. The results are shown in Table 1. The SBC framework achieves a mean absolute error (MAE) of 140.6. By utilizing integer-valued bins (E1), the MAE is remarkably reduced from 140.6 to 88.3–a significant 37.1% improvement. This outcome underscores the effectiveness of our approach in reducing ambiguity near the borders. By additionally adopting the noise reduction (E2) strategy, the MAE further decreases to 85.8. Performance improves even more when the combined loss (E3) is also used (i.e., by setting $\lambda$ in Eq. (3.1) to a positive value). Specifically, increasing the value of $\lambda$ from 0.00 to 1.00 results in improved performance, with the optimal result (77.9) achieved by EBCc with $\lambda=1.00$. However, an overemphasis on the count loss (e.g. $\lambda=2.00$) may deviate from the rational of blockwise classification, thus compromising the model’s generalizability.
We also verify that the performance improvement of EBC is agnostic to the choice of backbones We reimplement three other backbones (ResNet101 (He et al. 2016), MobileNetV2 (Sandler et al. 2018) and DenseNet201 (Huang et al. 2017)) used by the SBC paper (Liu et al. 2019), as well as two classic regression-based methods (CSRNet (Li, Zhang, and Chen 2018) and DMCount (Wang et al. 2020a)) within our EBC framework. The results in Table 2 show that our EBC framework consistently yields improved performance (up to 44.5%) across different backbones.
4.2 CLIP-EBC Compared with State-of-the-Art
@@
We compare CLIP-EBC with the latest crowd-counting methods. Table 3 presents the results on the ShanghaiTech A & B, UCF-QNRF, and NWPU-Crowd (Val) datasets, while Table  4 illustrates the results on the test split of the NWPU-Crowd database. Our CLIP-EBC models achieve comparable results with state-of-the-art approaches on the ShanghaiTech A & B and UCF-QNRF datasets. On the much larger benchmark database NWPU-Crowd, our CLIP-EBC model with the ViT-B/16 backbone achieves significant performance gains, surpassing all existing models with breakthrough results: an MAE of 36.6 and an RMSE of 81.7 on the validation split. On the test set, the ViT-B/16 backbone achieves an MAE of 61.3 and an RMSE of 278.4, outperforming all strong baselines. Moreover, the ViT-L/14 variant further reduces the errors, achieving an MAE of 58.2 and an RMSE of 268.5, setting new state-of-the-art performance.
4.3 Influence of Bin Granularity
Given that larger blocks can accommodate a wider range of possible count values, we investigate the effect of bin granularity on the performance of the EBC framework for block sizes $r=16$ and $r=32$. Table 5 shows the results on the ShanghaiTech A dataset using our CLIP-EBC (ViT-B/16) model. For both block sizes, the dynamic bin policy yields the best performance, benefitting from its ability to balance between reducing biases in average count values and increasing the sample sizes for undersampled bins. Additionally, in terms of MAE, the smaller block size $r=16$ produces better results, likely because it allows for better utilization of the spatial information present in the labels.
Granularity $\boldsymbol{r=16}$ $\boldsymbol{r=32}$
MAE RMSE MAE RMSE
Fine 54.32 90.31 54.95 89.14
Dynamic 53.80 88.66 54.07 89.11
Coarse 55.91 90.99 58.44 94.55
Table 5: Effects of the block size $r$ and the bin granularity. The experiments are conducted on the ShanghaiTech A dataset using our CLIP-EBC (ViT-B/16) model with different block sizes and bin policies. The results indicate that the dynamic bin strategy consistently outperforms the others.
Figure 4: Impact of different levels of noise reduction, controled by the factor $m$ in Eq. (3)-(5). Note that a smaller $m$ results in more annotations being treated as noise. The experiments are conducted on the ShanghaiTech A dataset with our CLIP-EBC (ViT-B/16) model. The results indicate that $m=4$ achieves the optimal effect.
4.4 Effects of Noise Control
A key component of our proposed EBC framework is the treatment of local count values greater than or equal to a threshold $m$ as noisy annotations. To mitigate the influence of these noisy labels, we group them into a single bin (Eq. (3)-(5)), thereby reducing their impact on the model’s performance. Here, we analyze how varying the threshold $m$ affects the performance of our model. We train our CLIP-EBC (ViT-B/16) model with different values of $m$ on ShanghaiTech A dataset (ranging from 1 to 9). Notably, the maximum local count value in the $8\times 8$ blocks of this dataset is 9, so setting $m=9$ implies that no labels are considered noisy. The results are illustrated in Fig. 4, which demonstrates that $m=4$ offers the best performance.
5 Conclusion
@@
In this paper, we bridged the gap between CLIP and crowd counting by reformulating the counting task as a blockwise classification problem. To this end, we proposed Enhanced Blockwise Classification (EBC), a novel, backbone-agnostic framework that improves upon standard blockwise classification through three key advancements: bin policy, noise control, and loss formulation. Within our EBC framework, we further proposed CLIP-EBC, the first fully CLIP-based crowd counting method, which successfully leverages CLIP for accurate crowd density estimation. CLIP-EBC classifies local count values into predefined integer-valued bins by comparing the similarity between local image features and corresponding text prompt features, followed by softmax to generate probability scores. Extensive experiments validated the effectivness of our EBC framework, and the comparison with existing methods on four benchmark datasets demonstrated the superior performance of our CLIP-EBC model. Future work will explore extending CLIP-EBC to address open-world object counting tasks.
Limitations
Although CLIP has been pre-trained on diverse image-text pairs, our work focuses solely on counting humans. In future work, we aim to extend this approach to general object counting.
Impact Statement
Like many other crowd counting methods, our models also analyze crowd images, which may raise privacy concerns, especially because their operation may create a sense of constant surveillance. To address these concerns, regulations and guidelines to govern the use of surveillance have been introduced in several countries (Fussey and Webster 2023).
Appendix A Impact of the Number of Learnable Tokens
We analyze the effect of varying the number of learnable tokens in our CLIP-EBC model with the ViT-B/16 backbone. The number of tokens is varied from 8 to 40, in increments of 4. The experimental results on the ShanghaiTech A dataset are shown in Fig. 5. As observed, increasing the number of learnable tokens from 8 to 32 results in a general improvement in performance, measured by both MAE and RMSE. The optimal performance is achieved at 32 tokens. However, beyond 32 tokens, the model’s performance begins to degrade, likely due to overfitting caused by the increased learnability.
Figure 5: Influence of the number of learnable tokens in our CLIP-EBC model with the ViT-B/16 backbone. The Experiments are conducted on the ShanghaiTech A dataset. The results show that setting the number of learnable tokens to 32 yields optimal performance, achieving the lowest values for both MAE and RMSE.
Appendix B More Experiments on Bin Granularity
We conducted further experiments to examine the impact of block size $r$ and bin granularity, this time using a VGG19-based encoder-decoder model (Ma et al. 2019; Wang et al. 2020a) as the backbone. This allows us to evaluate whether the dynamic binning strategy maintains its advantage even in non-CLIP models. All experiments were carried out on the UCF-QNRF dataset.
The results are presented in Table 6. Once again, the dynamic bin strategy proves to be more effective than both fine and coarse granularity, benefiting from its ability to strike an optimal balance between reducing biases in average count values and increasing sample sizes of undersampled bins. As for the impact of the block size $r$, under the MAE metric, performance slightly degrades when the block size increases to $r=32$. This decline can be attributed to the underutilization of spatial information as the block size grows. In particular, when the block size approaches the image size, the model is effectively trained to predict only the global count, and thus the spatial location information provided by point annotations no longer serves as supervision. In such cases, the task regresses into count-level weakly supervised crowd counting (Lei et al. 2021).
Granularity $\boldsymbol{r=16}$ $\boldsymbol{r=32}$
MAE RMSE MAE RMSE
Fine 76.31 131.74 76.94 130.87
Dynamic 75.90 130.48 76.06 127.72
Coarse 77.41 130.48 79.39 132.51
Table 6: Effects of the block size $r$ and the bin granularity. The experiments are conducted on the UCF-QNRF dataset using VGG19-based encoder-decoder model (Ma et al. 2019; Wang et al. 2020a) with different block sizes and bin policies. Consistent with the findings in Table 5, the dynamic bin strategy outperforms both fine and coarse binning approaches across both block sizes in terms of MAE and RMSE.
Appendix C Data Augmentation
In this section, we provide details of the data augmentation applied during the training in all our experiments. For each input image, we crop a region of size $su\times su$, where $s=224$ for ViT-based models, $s=448$ for all other models, and $u\sim\text{Uniform}[1,\,2]$. We then resize the cropped region to $s\times s$. This augmentation technique is designed to construct dense regions and incease the sample sizes of large bins, thereby alleviating undersampling issue caused by the long-tailed distribution of count values. Next, we apply a random horizontal flip with a probability of 0.5, followed by random color jittering with brightness=0.1, saturation=0.1, and hue=0.0. The hue factor is set to 0 to avoid NaN gradient issues during training, which were observed when positive values were used. Additionally, Gaussian blurring with a kernel size of 5 is applied to further enhance the robustness of the model Finally, we normalize the cropped region using the ImageNet mean and standard deviation.
Appendix D Visualization
We use our trained CLIP-EBC models to generate predicted density maps for four datasets: ShanghaiTech A & B, UCF-QNRF, and NWPU-Crowd. Specifically, CLIP-EBC (ResNet50) is applied to ShanghaiTech, CLIP-EBC (ViT-B/16) to UCF-QNRF, and CLIP-EBC (ViT-L/14) to NWPU-Crowd. A curated set of 19 representative images–six from ShanghaiTech, six from UCF-QNRF, and seven from NWPU-Crowd–are selected for visualization alongside their corresponding ground-truth and predicted density maps in Fig. 6.
  • Fig. 6(a) displays the results on ShanghaiTech, showing density maps predicted by CLIP-EBC (ResNet50).
  • Fig. 6(b) illustrates the results on UCF-QNRF using CLIP-EBC (ViT-B/16).
  • Fig. 6(c) presents the results on NWPU-Crowd with CLIP-EBC (ViT-L/14).
In each figure, the images progress from the sparsest to the most congested scenarios from left to right. These visualizations highlight the robustness of our models across a wide spectrum of crowd densities.
(a) Visualization on the ShanghaiTech (Zhang et al. 2016) dataset with our CLIP-EBC (ResNet-50) model.
(b) Visualization on the UCF-QNRF (Idrees et al. 2018) dataset with our CLIP-EBC (ViT-B/16) model.
(c) Visualization on the NWPU-Crowd (Wang et al. 2020b) dataset with our CLIP-EBC (ViT-L/14) model.
Figure 6: Visualization on the validation splits of different datasets using different CLIP-EBC models. Top to bottom rows: original images, the Gaussian-smoothed ground-truth density maps, and the density maps predicted by CLIP-EBC. Columns from the left to right represent scenarios of increasing density.

### v3

@@ III-B The Structure of CLIP-EBC
Next, the resulting $n$ text prompts are tokenized using CLIP’s tokenizer. We then input the tokenized text into the original CLIP text encoder, with its parameters frozen during training. This process yields text features $\boldsymbol{F}^{\text{txt}}$ with dimensions $d\times n$. The probability map $\boldsymbol{P}^{*}$ can be computed from the image feature map $\boldsymbol{F}^{\text{img}}$ and the text features $\boldsymbol{F}^{\text{txt}}$. First, we calculate the cosine similarity between the image feature vector $\boldsymbol{F}^{\text{img}}_{:,\,i,\,j}$ and the $n$ extracted text embeddings. We then normalize these similarities using softmax to obtain $\boldsymbol{P}^{*}_{:,\,i,\,j}$, which denotes the probability scores across the $n$ bins for block $(i,\,j)$. To obtain the predicted density map $\boldsymbol{Y}^{*}$, we use Eq. (2). We train CLIP-EBC with our loss function in Eq. (III-A4).
IV Experiments
@@
Implementation Details. We set the block size $r=8$ (unless explicitly specified), which is commonly used in many works [5, 8, 6]. For this block size, we set $m=4$. We use the Adam optimizer [17] to train all our models with an initial learning rate of $41e-$ 4, which is adjusted through a cosine annealing schedule. The batch size is fixed at $8$ for all datasets. For our ViT-based CLIP-EBC model, we apply visual prompt tuning [18] by prepending 32 learnable tokens to each layer, rather than fine-tuning the entire backbone. Following existing methods, we use the mean absolute error (MAE) and root mean square error (RMSE) to evaluate our models.
IV-A Effectiveness of EBC
We first demonstrate the effectiveness of our EBC framework using VGG16 [19] as the backbone, trained and evaluated on the UCF-QNRF dataset. The results are shown in Table I. The SBC framework (i.e., the scenario where no enhancement is used) achieves a mean absolute error (MAE) of 140.6. By utilizing integer-valued bins (E1), the MAE is remarkably reduced from 140.6 to 88.3–a significant 37.1% improvement. This outcome underscores the effectiveness of our approach in reducing ambiguity near the borders. By additionally adopting noise reduction (E2), the MAE further decreases to 85.8. Performance improves even more when the combined loss (E3) is also used. Specifically, increasing the value of $\lambda$ from 0.00 to 1.00 results in improved performance, with the optimal result (77.9) achieved with $\lambda=1.00$.
We also verify that the performance improvement of EBC is agnostic to the choice of backbones. We re-implement three other backbones (ResNet101 [20], MobileNetV2 [21] and DenseNet201 [22]) used by the SBC paper [8], as well as two classic regression-based methods (CSRNet [5] and DMCount [6]) within our EBC framework. The results in Table II show that our EBC framework consistently yields improved performance (up to 44.5%) across different backbones.
IV-B CLIP-EBC Compared with State-of-the-Art
@@
P2PNet [13] 72.6 331.6
CLTR [14] 74.4 333.8
STEERER [15] 63.7 309.8
CrowdHat [23] 68.7 296.9
PET [24] 74.4 328.5
@@
TABLE IV: Comparison of our model CLIP-EBC with state-of-the-art methods on NWPU-Crowd (Test) [11]. CLIP-EBC achieves both the lowest MAE and RMSE on this large-scale benchmark database, surpassing all strong baselines.
@@
We compare CLIP-EBC with the latest crowd-counting methods. Table III presents the results on the ShanghaiTech A & B, UCF-QNRF, and NWPU-Crowd (Val) datasets, while Table  IV illustrates the results on the test split of the NWPU-Crowd database. Our CLIP-EBC models achieve comparable results with state-of-the-art approaches on the ShanghaiTech A & B and UCF-QNRF datasets. On the much larger benchmark database NWPU-Crowd, our CLIP-EBC model with the ViT-B/16 backbone achieves significant performance gains, surpassing all existing models with breakthrough results: an MAE of 36.6 and an RMSE of 81.7 on the validation split. On the test set, the ViT-B/16 backbone achieves an MAE of 61.3 and an RMSE of 278.4, outperforming all strong baselines. Moreover, the ViT-L/14 variant further reduces the errors, achieving an MAE of 58.2 and an RMSE of 268.5, setting new state-of-the-art performance. Fig. 4 provides visualized results on the validation split using CLIP-EBC with the ViT-L/14 backbone.
Figure 4: Visualization of density maps predicted by CLIP-EBC.
V Conclusion
@@
In this paper, we bridged the gap between CLIP and crowd counting by reformulating the counting task as a blockwise classification problem. To this end, we proposed Enhanced Blockwise Classification (EBC), a novel, backbone-agnostic framework that improves upon standard blockwise classification through three key advancements: quantization, noise reduction, and loss formulation. Within our EBC framework, we further proposed CLIP-EBC, the first fully CLIP-based crowd counting method, which successfully leverages CLIP for accurate crowd density estimation. CLIP-EBC classifies local count values into predefined integer-valued bins by comparing the similarity between local image features and text prompt features. Extensive experiments validated the effectiveness of our EBC framework, and the comparison with existing methods on four benchmark datasets demonstrated the superior performance of our CLIP-EBC model.